Circular No. 6549

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1997 C1 Tom Gehrels, Lunar and Planetary Laboratory, reports his discovery

of a comet by eye in the course of the Spacewatch program. Available CCD precise positions:

1997 UT  α_{2000}  δ_{2000}  m_{1} Observer
Feb. 1.36102 9^{h}11^{m}33.^{s}00 +12^{o}17′35′′.0 17.9 Gehrels
1.38901 9 11 32.08 +12 17 38.9  _{”}
1.41641 9 11 31.13 +12 17 43.0  _{”}
2.01610 9 11 11.05 +12 19 13.4 17.9 Tichá
2.01816 9 11 11.00 +12 19 13.4  _{”}
2.02000 9 11 10.99 +12 19 13.8  _{”}
2.02227 9 11 10.87 +12 19 14.6  _{”}
2.02534 9 11 10.71 +12 19 14.8  _{”}

T. Gehrels (Kitt Peak). 0.9-m Spacewatch telescope. Condensation of

diameter 7′′. Fuzzy tail, 0′.5 long in p.a. 280^{o}. J. Tichá and M. Tichý (Kleť). 0.57-m f/5.2 reflector. Object diffuse.

SUPERNOVAE 1997U AND 1997V IN ANONYMOUS GALAXIES B. Schmidt reports the discovery by the Mount Stromlo Abell Cluster

Supernova Search Team (cf. IAUC 6418) of two apparent supernovae near Abell 423 on CCD images taken by S. Sabine with the Mount Stromlo 1.27m telescope (+ Macho Camera). SN 1997U is located at α = 3^{h}10^{m}46^{s}.25, δ = −12^{o}16′04′′.5 (equinox 2000.0), which is 0′′.01 east and 0′′.45 north of the nucleus of the host galaxy. SN 1997U is visible at m_{R} ∼ 21.5 on V and R CCD images taken on Jan. 2, and it was also detected on similar frames taken on Jan. 5 and 8. SN 1997V is located at α = 3^{h}11^{m}43^{s}.82, δ = −12^{o}02′11′′.0 (equinox 2000.0), which is 6′′.0 east and 0′′.1 north of the nucleus of the host galaxy. SN 1997V is visible at m_{R} ∼ 21.0 on V and R CCD images taken on Jan. 8, and is also present, but 0.5 mag fainter, on similar frames taken on Jan. 5.

Neither SN 1997U nor SN 1997V was detected on frames taken on 1996 Dec. 18. Finding charts are available at http://msowww.anu.edu.au/

~reiss/Abell SNSearch/.

1997 February 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green